There are 150 rectangular tiles that are 2 inches by 3 inches. They are all used to cover a square region with no gaps or overlaps between the tiles. No tiles are cut or broken in the process. What is the perimeter of the square that is covered?
a. It is impossible to cover a square with these tiles.
b. The perimeter is 600 inches.
c. The perimeter is 200 inches.
d. The perimeter is 300 inches.
e. The perimeter is 120 inches.

Answer by CharlesG2(834)

There are 150 rectangular tiles that are 2 inches by 3 inches. They are all used to cover a square region with no gaps or overlaps between the tiles. No tiles are cut or broken in the process. What is the perimeter of the square that is covered?
a. It is impossible to cover a square with these tiles.
b. The perimeter is 600 inches.
c. The perimeter is 200 inches.
d. The perimeter is 300 inches.
e. The perimeter is 120 inches.

150 tiles, 2 inches by 3 inches
150 * 2 * 3 = 150 tiles * 6 square inches = 900 square inches to be covered by the tiles
a square with area 900 square inches needs to have 30 inches per side, since 30*30=900
this square would have a perimeter of 120 inches, since 4*30=120

the tiles would be arranged 10 tiles along one side (10*3=30) by 15 tiles along the other way (15*2=30)

the answer is e.